# Windowed MAPF with Completeness Guarantees

### v1

@@ 5.2 Tough Congested Scenarios
The previous section shows how SS-CBS can outperform windowed ECBS in standard benchmark scenarios. We were additionally interested in SS-CBS’s performance in extremely congested scenarios and thus evaluated it on small maps with high congestion from Okumura (2022).
Table 1 shows the average success from 20 seeds with a timeout of 1 minute, with the number $N$ in parenthesis denoting using the first $N$ agents in a scenario. We see how windowed ECBS with window sizes $1,2,4,8,16$ fails completely. Additionally, we compared it to EECBS (Explicit Estimation high-level search) with all CBS improvements, including bypass, symmetry reasoning, prioritized conflicts, and full horizon planning. Despite these enhancements, EECBS still struggles due to high congestion.
SS-CBS can solve nearly all problems with high accuracy. It’s worth emphasizing that SS-CBS uses no CBS improvements, relying solely on basic focal search with $w_{so}=1$ and planning just a single step at a time. Despite this, it outperforms both windowed ECBS and EECBS baselines. Additional results in the appendix further compare SS-CBS to EECBS, demonstrating that SS-CBS is at least an order of magnitude faster on solved instances.
SS-CBS’s performance demonstrates two key points. First, its high success rate on these challenging, congested maps shows how heuristic penalties can effectively guide SS-CBS to bypass complex congestion. Second, EECBS struggles due to the high number of conflicts in these scenarios. SS-CBS’s success highlights how iterative single-step planning with heuristic updates can be an effective approach for resolving difficult congestion issues. All existing windowed methods produce congestion due to their myopic planning. Thus, SS-CBS improving performance with windowed planning in severe congestion is remarkable.
@@ 6 Conclusion and Future Work
Existing MAPF works have focused on designing methods to solve full-horizon planning. When faced with shorter deadlines, all current methods take the full horizon MAPF methods and simply reduce the planning horizon to a smaller window. This has been shown to cause deadlock/livelock/insurmountable congestion due to the limited planning horizon. We introduce WinC-MAPF, the first framework that enables theoretical completeness with windowed MAPF solvers. In particular, we show that using windowed “Action Generator” that incorporates heuristic penalties, identifies agent groups, and optimally minimizes $c(\mathcal{C},\mathcal{C}^{\prime})+h(\mathcal{C}^{\prime})$ is complete. Following this framework, we designed SS-CBS which uniquely introduces “heuristic conflicts” to successfully incorporate heuristic penalties and return the optimal next step action. We experimentally validate how our theoretically complete method actually translates to real performance benefits with SS-CBS consistently outperforming windowed ECBS across a variety of windows and maps.
We are very excited about future work that can relax the limitations of our framework. First, the most useful extension is to incorporate and prove completeness for bounded suboptimal AGs within our framework. This can allow more solvers such as PIBT, MAPF-LNS2 (Li et al. 2022), W-EECBS (Veerapaneni, Kusnur, and Likhachev 2023), or even using a learnt neural network policy. Second, one obvious extension is to generalize SS-CBS to work on longer horizons. This also opens the door to using more sophisticated real-time heuristic update methods. Third, future work could try to relax the perfect backward Dijkstra single-agent heuristic assumption and learn individual heuristics online.
@@ Appendix A Quick Summary
Readers new to MAPF or (E)CBS are recommended to read ECBS (Barer et al. 2014). Readers unfamiliar with windowed MAPF solvers should read Rolling-Horizon Collision Resolution (Li et al. 2020). Readers new to Real-Time Heuristic Search can take a look at LRTA* in Korf (1990).
@@
The majority of MAPF methods which find entire collision free paths to goal can take a long time (e.g. $>10$ seconds). Real-world practitioners cannot wait this long. Thus to reduce planning time, existing works only reason about collisions within a fixed time window/horizon, where the window $W$ is much smaller than the entire solution path.
A key issues with these windowed approaches is that their myopic planning results in deadlock or livelock if their window is too small. Table 1 shows examples where windowed MAPF solvers fail in congestion which requires longer horizon planning. More broadly, all existing windowed MAPF solvers regardless of window size lack theoretical completeness and several windowed works have explicitly cited deadlock as a key issue in their experiments (Li et al. 2020; Okumura et al. 2022; Jiang et al. 2024).
@@
1. Windowed Complete MAPF (WinC-MAPF) Framework: We develop the first general framework that enables creating windowed MAPF solvers that have completeness guarantees. Our first insight is that we can leverage the single-agent Real-Time Heuristic Search perspective which uses limited horizon/windowed planning but maintains completeness by updating the heuristics of visited states. Our second insight is to leverage the semi-independence of agents in MAPF and only computing the heuristic updates in-respect to groups of coupled agents rather than the entire joint configuration space.
Formally, we define a windowed MAPF “Action Generator” (AG) as a search method that given a window $W$ and a current configuration $\mathcal{C}$, finds a $\mathcal{C}^{\prime}$ (and associated actions) that is a valid neighboring configurations within $W$ timesteps. Additionally, the AG needs to determine groups of coupled agent groups as defined in Section 3.2. We prove in the next section how an optimal windowed AG that computes $\operatorname*{arg\,min}c(\mathcal{C},\mathcal{C}^{\prime})+h(\mathcal{C}^{\prime})$ is complete under standard conditions which regular MAPF holds. It is likely possible that suboptimal AG’s, e.g. ones that do not exactly find the argmin but some bounded suboptimal AG, will be complete but we were unable to prove that using standard techniques and is left for future work.
2. Single-Step CBS (SS-CBS): We develop SS-CBS which is an instantiation of CBS following our WinC-MAPF framework. SS-CBS finds the optimal next step (so window $W=1$) given heuristic updates. We show that naively incorporating heuristic updates (which we redefine in respect to heuristic “penalties”) by adding it into the high-level or low-level CBS search is incorrect. SS-CBS’s innovation is to incorporate heuristic penalties by introducing a new “heuristic conflict” and constraint that defers the addition of the heuristic penalty and enables all agents to replan to avoid a penalty. We additionally show how SS-CBS can easily determine coupled agent groups by merging pairs of conflicting agents.
Experimentally, Figure 3 shows how SS-CBS has a higher success rate and agent scalability on standard MAPF benchmark maps when compared to windowed ECBS across a windows $w=\{1,2,4,8,16\}$. We highlight that SS-CBS plans only one single-step (i.e. extremely myopic planning) but is able to handle congestion/deadlock better. Additionally, we investigate the performance of these windowed methods tough small scenarios (Table 1, A1). In these scenarios, we find that windowed ECBS methods uniformly fail, and that even EECBS with all optimizations struggles. SS-CBS is able to almost perfectly solve these instances. All existing windowed MAPF methods struggle in congestion, so SS-CBS’s superior performance in these congested scenario highlights the power of the WinC-MAPF framework.
@@
2. SS-CBS requires a perfect single-agent heuristic. Although several other MAPF methods requires this (e.g. PIBT/LaCAM (Veerapaneni et al. 2024)), this is not generally required for CBS and future work could relax this.
@@
4. As mentioned earlier, our WinC-MAPF currently proves completeness with an optimal AG. We conjecture that future work can likely modify and prove that bounded suboptimal AG’s can be complete within our framework.
Appendix B Proving Completeness of WinC-MAPF with Agent Group Heuristic Penalties
A key idea of WinC-MAPF is to compute heuristic penalties on groups of agents rather than the entire joint configuration. Without this, i.e. if we just computing heuristic penalties on the full joint configuration space, we get completeness by directly applying the single-agent proof. We thus first define agent groups. We then roughly restate the standard proof used in the original real-time search LRTA* paper (Korf 1990) to provide context on the differences that our WinC-MAPF framework has in respect to proving completeness. We cannot directly use LRTA*’s proof. Instead, we can use the fact that this proof shows how (1) proving that an algorithm doesn’t cycle infinitely proves completeness and (2) how heuristic values with cycles with an update of $h^{\prime}(s_{i})\leftarrow c(s_{i},s_{i+1})+h(s_{i+1})$ results in infinitely large heuristic values. We then show how our heuristic values are admissible, which combined with (1) and (2) with optimal windowed AGs result in our framework being complete.
B.1 Agent Groups
The general definition of a group of coupled agents is a set of agents where for each agent $R_{i}$, there exists an $R_{j}$ that either blocks its path or vice-versa. From an abstract perspective, we can compute this by iterating through each agent which is not on its optimal single-agent path and storing the ids of the other agents which prevented it from picking a better path (there must exist at least one other agent otherwise the agent could have gone on its optimal path). This builds a dependency graph where agents that share an edge $R_{i}\rightarrow R_{j}$ denote $R_{i}$ blocked by $R_{j}$.
We can then find all the disjoint connect components in this dependency graph (e.g. via a DFS) where each disjoint connected component depicts a group of agents that are blocking each other. Instead of doing $h(\mathcal{C})\leftarrow c(\mathcal{C},\mathcal{C}^{\prime})+h(\mathcal{C}^{\prime})$, we do $h(\mathcal{C}_{Gr})\leftarrow c(\mathcal{C}_{Gr},\mathcal{C}_{Gr}^{\prime})+h(\mathcal{C}_{Gr}^{\prime})$ for each group of agents $Gr$ at configuration $\mathcal{C}_{Gr}$.
B.2 Standard Single-Agent Real-Time Search Completeness Proof
LRTA* (Korf 1990) is a single agent algorithm where at each timestep, the agent picks $\operatorname*{arg\,min}_{s^{\prime}}c(s,s^{\prime})+h(s^{\prime})$ over successor states $s^{\prime}$ and updates the heuristic $h^{\prime}(s)\leftarrow c(s,s^{\prime})+h(s^{\prime})$. We summarize its proof of completeness.
@@ B.2 Standard Single-Agent Real-Time Search Completeness Proof
In a finite problem space with positive edge costs and finite heuristic values, in which a goal state is reachable from every state, LRTA* will find a solution.
@@
In a finite state space, if LRTA* does not reach the goal, there must exist a finite cycle that LRTA* is stuck in (otherwise it will visit new states and eventually the goal).
@@
LRTA* is therefore guaranteed to eventually leave any finite cycle. Since a finite problem space has a finite set of cycles that do not include the goal, LRTA* at worst will explore all of these and eventually exit them and reach the goal. ∎
B.3 WinC-MAPF Proof
@@
A key idea of WinC-MAPF is to compute heuristic penalties on groups of agents rather than the entire joint configuration. Note that without this, i.e. if we just computing heuristic penalties on the full joint configuration space, we get completeness by directly applying the single-agent proof. Thus, using heuristic penalties on groups complicates the proof of completeness as when computing/creating a heuristic penalty, we are increase the heuristic values of joint-states not in the cycle. We unfortunately cannot use this proof.
@@
Given a finite undirected search space and: (1) the initial heuristic is admissible, (2) our AG picks $\mathcal{C}^{\prime}=\operatorname*{arg\,min}c(\mathcal{C},\mathcal{C}^{\prime})+h(\mathcal{C}^{\prime})$ and identifies coupled agents groups, then WinC-MAPF with its grouped update equation will always have admissible heuristic values for all $\mathcal{C}$.
We calculate the joint heuristic value by summing up mutually exclusive group heuristic penalties as described in Section C.1. If we can prove that the group heuristic values are admissible, then the sums of disjoint grouped heuristic values is also admissible as groups interacting can only increase the true cost-to-go. Thus we prove the following lemma.
@@
Given a finite undirected search space and: (1) the initial heuristic is admissible, (2) our AG picks $\mathcal{C}^{\prime}=\operatorname*{arg\,min}c(\mathcal{C},\mathcal{C}^{\prime})+h(\mathcal{C}^{\prime})$ and identifies coupled agents groups, then WinC-MAPF with its grouped update equation will always have admissible heuristic values for all group configurations $\mathcal{C}_{Gr}$.
@@
We prove this via induction. Our inductive hypothesis is that we have $h(\mathcal{C}_{Gr})\leq h^{*}(\mathcal{C}_{Gr})$ across all grouped agent $Gr$ configurations $\mathcal{C}_{Gr}$ and iterations of running our Windowed MAPF framework.
@@
Inductive Step: We assume that at some timestep $T$ at $\mathcal{C}$, all the heuristic values are admissible. The AG picks a $\mathcal{C}^{\prime}$ that minimizes $c(\mathcal{C},\mathcal{C}^{\prime})+h(\mathcal{C}^{\prime})$. Then for each coupled group of agents, we update the group’s joint-configuration (via our heuristic penalty) to $h(\mathcal{C}_{Gr})\leftarrow c(\mathcal{C}_{Gr},\mathcal{C}_{Gr}^{\prime})+h(\mathcal{C}_{Gr}^{\prime})$.
Suppose that there is a group of agents $GrA$ at $\mathcal{C}_{GrA}$ whose heuristic value becomes inadmissible. This means that the chosen $\mathcal{C}_{GrA}^{\prime}$ was not optimal and there existed a different $\mathcal{C}_{GrA}^{*}$ that was optimal. However, since we are running on optimal AG, we would have replaced $\mathcal{C}_{GrA}^{\prime}$ with $\mathcal{C}_{GrA}^{*}$. If this did not interact with any other agents, this would reduce our overall cost without changing any other agents positions/actions and contradict the optimal AG. If it did interact with other agents, then $GrA$ would need to be larger and include these other agents, violating our definition of groups. Thus, neither of these are possible. This means that our heuristic value stays admissible for all groups after updating. ∎
Appendix C SS-CBS
C.1 Subtleties
Determining Which HPs to Apply
As discussed earlier, give a CT node we need to detect heuristic conflicts. This can be done by going through all the heuristic penalties and checking if their agent locations match with the CT’s configuration.

### v2

@@ 5.2 Tough Congested Scenarios
The previous section shows how SS-CBS can outperform wCBS in standard benchmark scenarios. We were additionally interested in SS-CBS’s performance in extremely congested scenarios and thus evaluated it on small maps with high congestion from Okumura (2022).
Table 1 shows the average success from 20 seeds with a timeout of 1 minute, with the number $N$ in parenthesis denoting using the first $N$ agents in a scenario. We see how windowed CBS with $W=1,2,4,8,16$ fails completely. Additionally, CBS with all CBS improvements, including bypass, symmetry reasoning, prioritized conflicts, and full horizon planning (CBS+) struggles.
Table 2 shows the average runtime in milliseconds. On solved instances, SS-CBS is much faster compared to EECBS with a suboptimality $w_{so}$ of 1.5 and CBS+. Limitation: However, the drawback of using SS-CBS here is that its solution cost is horrible. It is about 7500 for Tunnel (4), 11000 for Loopchain (7), and 450 for Connector (6).
SS-CBS CBS+ EECBS $w_{so}=1.5$
Instance Success Time Success Time Success Time
Tunnel (3) 1 85 0.9 43,423 0.80 3,768
Tunnel (4) 1 2,713 0 - 0 -
Loopchain (6) 1 15,393 0 - 0 -
Loopchain (7) 0.95 18,268 0 - 0 -
Connector (5) 1 15 0 - 0.95 146
Connector (6) 1 43 0 - 0.90 1,116
Table 2: SS-CBS with one-step planning can solve problems that even CBS and EECBS with full horizon planning and all optimizations cannot. This highlights SS-CBS’s and the broader WinC-MAPF framework’s ability to effectively use grouped heuristic penalties to resolve congestion. Time (ms) is averaged over successful instances.
SS-CBS’s performance demonstrates two key points. First, its high success rate on these challenging, congested maps shows how heuristic penalties can effectively guide SS-CBS to bypass complex congestion. Second, full horizon CBS+ struggles due to the high number of conflicts in these scenarios. SS-CBS’s success highlights how iterative single-step planning with heuristic updates can be an effective approach for resolving difficult congestion issues. All existing windowed methods produce congestion due to their myopic planning. Thus, SS-CBS improved success rate with windowed planning in severe congestion is remarkable.
@@ 6 Conclusion and Future Work
Existing MAPF works have focused on designing methods to solve full-horizon planning. When faced with shorter deadlines, all current methods take the full horizon MAPF methods and simply reduce the planning horizon to a smaller window. This has been shown to cause deadlock/livelock/insurmountable congestion due to the limited planning horizon. We introduce WinC-MAPF, the first framework that enables theoretical completeness with windowed MAPF solvers. In particular, we show that using windowed “Action Generator” that incorporates heuristic penalties, identifies disjoint agent groups, and optimally minimizes $c(\mathcal{C},\mathcal{C}^{W})+h(\mathcal{C}^{W})$ is complete. Following this framework, we designed SS-CBS which uniquely introduces “heuristic conflicts” to successfully incorporate heuristic penalties and return the optimal next step action. We experimentally validate how our theoretically complete method actually translates to real performance benefits with SS-CBS consistently outperforming windowed CBS across a variety of windows and maps.
We are very excited about future work that can relax the limitations of our framework. First, the most useful extension is to incorporate and prove completeness for bounded or even abitrarily suboptimal AGs within our framework (we are limited to optimal AGs). This can allow more solvers such as PIBT, MAPF-LNS2 (Li et al. 2022), W-EECBS (Veerapaneni, Kusnur, and Likhachev 2023), or even using a learnt neural network policy. Second, one obvious extension is to generalize SS-CBS to work on longer horizons. This also opens the door to using more sophisticated real-time heuristic update methods. Third, future work could try to relax the perfect backward Dijkstra single-agent heuristic assumption and learn individual heuristics online.
@@
Acknowledgments
The research was supported by the National Science Foundation under grant #2328671, by the National Science Foundation Graduate Research Fellowship Program under grant #DGE2140739, and a gift from Amazon. The views and conclusions in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the sponsoring organizations, agencies, or the U.S. government.
@@ Appendix A Quick Summary
Readers new to MAPF or CBS are recommended to read CBS (Sharon et al. 2015). Readers unfamiliar with windowed MAPF solvers should read Rolling-Horizon Collision Resolution (Li et al. 2020). Readers new to Real-Time Heuristic Search can take a look at LRTA* (Korf 1990).
@@
The majority of MAPF methods which find entire collision-free paths to goal can take a long time (e.g. $>10$ seconds). Real-world practitioners cannot wait this long. Thus, to reduce planning time, existing works only reason about collisions within a fixed time window/horizon, where the window $W$ is much smaller than the length of the entire solution path.
A key issue with these windowed approaches is that their myopic planning results in deadlock or livelock if their window is too small. Table 1 shows examples where windowed MAPF solvers fail in congestion which requires longer horizon planning. More broadly, all existing windowed MAPF solvers regardless of window size lack theoretical completeness, and several windowed works have explicitly cited deadlock as a key issue in their experiments (Li et al. 2020; Okumura et al. 2022; Jiang et al. 2024).
@@
1. Windowed Complete MAPF (WinC-MAPF) Framework: We develop the first general framework that enables creating windowed MAPF solvers that have completeness guarantees. Our first insight is that we can leverage the single-agent Real-Time Heuristic Search perspective which uses limited horizon/windowed planning but maintains completeness by updating the heuristics of visited states. This, however, does not lead to a practical algorithm as it requires exploring the entire joint configurations to get out of deadlocks/livelocks. Our second insight is to leverage the semi-independence of agents in MAPF and only compute the heuristic updates in respect to groups of coupled agents rather than the entire joint configuration space.
Formally, we define a windowed MAPF “Action Generator” (AG) as a search method that, given a window $W$ and a current configuration $\mathcal{C}$, finds a $\mathcal{C}^{W}$ (and associated actions) that is a valid neighboring configuration within $W$ timesteps. Additionally, the AG needs to determine groups of coupled agent groups as defined in Section 3.2. We prove in the next section how an optimal windowed AG that computes $\operatorname*{arg\,min}c(\mathcal{C},\mathcal{C}^{W})+h(\mathcal{C}^{W})$ is complete under regular MAPF conditions.
2. Single-Step CBS (SS-CBS): We develop SS-CBS which is an instantiation of an AG that can be used within our WinC-MAPF framework. SS-CBS finds the optimal next step (so window $W=1$) given heuristic updates. We show that naively incorporating heuristic updates (which we redefine as heuristic “penalties”) by adding it into the high-level or low-level CBS search is incorrect. SS-CBS’s innovation is to incorporate heuristic penalties by introducing a new “heuristic conflict” and constraint that defers the addition of the heuristic penalty and enables all agents to replan to avoid a penalty. We additionally show how SS-CBS can easily determine coupled agent groups by merging pairs of conflicting agents.
Experimentally, Figure 3 shows how SS-CBS has a higher success rate and agent scalability on standard MAPF benchmark maps compared to windowed CBS across windows $W=\{1,2,4,8,16\}$. We highlight that SS-CBS plans only one single step (i.e., extremely myopic planning) but is able to handle congestion/deadlock better. Additionally, we investigate the performance of these windowed methods in tough small scenarios (Tables 1 and 2). In these scenarios, we find that windowed CBS methods uniformly fail, and that even CBS with all optimizations struggles. SS-CBS is able to almost perfectly solve these instances. All existing windowed MAPF methods struggle in congestion, so SS-CBS’s superior performance in these congested scenarios highlights the power of the WinC-MAPF framework.
@@
2. SS-CBS requires a perfect single-agent heuristic. Although most MAPF methods assume this, this is not generally required for CBS and future work could relax this.
@@
4. Our WinC-MAPF currently proves completeness with an optimal AG. We conjecture that future work can likely modify and prove that bounded suboptimal AG’s can be complete within our framework.
Appendix B Proving Completeness of WinC-MAPF with Group Heuristic Penalties
A key idea of WinC-MAPF is to compute heuristic penalties (HPs) on groups of agents rather than the entire configuration. We first roughly restate the standard single-agent proof used in the original real-time search LRTA* paper (Korf 1990) to provide context on the differences that our WinC-MAPF framework has in respect to proving completeness. We cannot directly use LRTA*’s proof due to our use of agent group HPs. Instead, we can use the fact that this proof shows how (1) proving that an algorithm does not cycle infinitely proves completeness and (2) how heuristic values with cycles with an update of $h^{\prime}(s^{i})\leftarrow c(s^{i},s^{i+1})+h(s^{i+1})$ results in infinitely large heuristic values. We then show how our heuristic values are admissible, which, combined with (1) and (2) with optimal windowed AGs, result in our framework being complete.
B.1 Standard Single-Agent Real-Time Search Completeness Proof
LRTA* (Korf 1990) is a single-agent algorithm where at each timestep, the agent picks $\operatorname*{arg\,min}_{s^{\prime}}c(s,s^{\prime})+h(s^{\prime})$ over successor states $s^{\prime}$ and updates the heuristic $h^{\prime}(s)\leftarrow c(s,s^{\prime})+h(s^{\prime})$. We summarize its proof of completeness.
@@ B.1 Standard Single-Agent Real-Time Search Completeness Proof
In a finite bidirectional graph with positive edge costs and finite heuristic values, in which a goal state is reachable from every state, LRTA* will find a solution.
@@
In a finite bidirectional graph, if LRTA* does not reach the goal, there must exist a finite cycle that LRTA* is stuck in (otherwise it will visit new states and eventually the goal).
@@
Thus, LRTA* is guaranteed to eventually leave any finite cycle. Since a finite bidirectional graph has a finite set of cycles without the goal, LRTA* at worst will explore all of these and eventually exit them and reach the goal. ∎
Subtleties
This proof for completeness assumes that $h^{\prime}(s_{i})\leftarrow c(s_{i},s_{i+1})+h(s_{i+1})\geq h(s_{i})$. This holds when $h$ is consistent but may not hold when $h$ is admissible. Thus for arbitrary $h$ we should update via $h^{\prime}(s_{i})\leftarrow\max(h(s_{i}),c(s_{i},s_{i+1})+h(s_{i+1}))$ (Eq. 1) and then reuse the same completeness proof.
This proof also assumes that $s_{i-1}$ has a different state that can be picked, i.e. it has more than one neighbor. If there exists a path in the cycle to the goal, then there must be some $s_{i-1}$ where this holds and we apply this logic for that $s_{i-1}$.
Lastly, this proof was shown for one-step planning but can be directly applied to planning for $W$ steps. It is also independent of how many steps along the $W$ length partial path the agent chooses to move (as long as it moves at least one step along the path).
B.2 WinC-MAPF Proof
@@
A key idea of WinC-MAPF is to compute heuristic penalties on the configuration of disjoint agent groups rather than the entire configuration. Note that without this, i.e., if we just compute heuristic penalties on the full configuration, we get completeness by directly applying the single-agent proof. Thus, using group HPs complicates the proof of completeness as when computing/creating a heuristic penalty, we increase the heuristic values of configurations not in the cycle.
@@
Given a finite bidirectional graph and: (1) an initial Backward Dijkstra heuristic, (2) our AG picks $\operatorname*{arg\,min}_{\mathcal{C}^{W}}c(\mathcal{C},\mathcal{C}^{W})+h(\mathcal{C}^{W})$ and identifies disjoint agents groups, then WinC-MAPF with its update equation (Eq. 1) applied on group configurations will always have admissible heuristic values for all $\mathcal{C}$.
We calculate the joint heuristic value by summing up mutually exclusive group HPs as described in Section C.1. If we can prove that each group HPs values are admissible, then the sums of disjoint grouped heuristic values is also admissible as groups interacting can only increase the true cost-to-go. Thus we prove the following lemma.
@@
Given a finite bidirectional graph and: (1) an initial Backward Dijkstra heuristic, (2) our AG picks $\operatorname*{arg\,min}_{\mathcal{C}^{W}}c(\mathcal{C},\mathcal{C}^{W})+h(\mathcal{C}^{W})$ and identifies disjoint agents groups, then WinC-MAPF with its update equation (Eq. 1) applied on group configurations will always have admissible heuristic values for all group configurations $\mathcal{C}_{Gr}$.
@@
We prove this via induction. Our inductive hypothesis is that we have $h(\mathcal{C}_{Gr})\leq h^{*}(\mathcal{C}_{Gr})$ across all group $Gr$ configurations $\mathcal{C}_{Gr}$ and iterations of running our Windowed MAPF framework.
@@
Inductive Step: We assume that at some timestep $T$ at $\mathcal{C}$, all the heuristic values are admissible. The AG picks a $\mathcal{C}^{W}$ that minimizes $c(\mathcal{C},\mathcal{C}^{W})+h(\mathcal{C}^{W})$. Then for each group, we update its configuration (via our heuristic penalty) to $h(\mathcal{C}_{Gr})\leftarrow U(\mathcal{C}_{Gr},\mathcal{C}_{Gr}^{W})+h(\mathcal{C}_{Gr}W)$.
Suppose that there is a group $GrA$ at $\mathcal{C}_{GrA}$ whose heuristic value became inadmissible after applying $U$. This can only occur when the chosen $\mathcal{C}_{GrA}^{W}$ was not optimal and there must exist a different $\mathcal{C}_{GrA}^{*}$ where $U(\mathcal{C}_{GrA},\mathcal{C}_{GrA}^{*})$ would be admissible. However, since we are running on optimal AG, we would have replaced $\mathcal{C}_{GrA}^{\prime}$ with $\mathcal{C}_{GrA}^{*}$. If this did not interact with any other agents, this would reduce our overall cost without changing any other agents’ locations/actions and contradict the optimal AG. If it did interact with other agents, then $GrA$ would need to be larger and include these other agents, violating our definition of disjoint agent groups. Thus, neither of these are possible. This means that our heuristic value stays admissible for all disjoint agent groups after updating. ∎
Appendix C SS-CBS Subtleties
C.1 Determining Which HPs to Apply
As discussed earlier, give a CT node we need to detect heuristic conflicts. This can be done by going through all the HPs and checking if their agent locations match with the CT’s configuration.